OpenStudy (anonymous):

A line passes through the points (-6, 7) and (x2, y2), and has a slope of -1/12. If point (x2, y2) is located in quadrant I, find y2

OpenStudy (anonymous):

use m=(y1-y2)/(x1-x2) to obtain y=-x/12+13/2 so 13/2 is the y intercept and 78 is the x intercept. y exists in quadrant one from 0 to 13/2 Domain: (0,13.5)

OpenStudy (anonymous):

Range I mean, not domain